# Data-dependent and Oracle Bounds on Forgetting in Continual Learning

### v2

@@ Appendix A Appendix - proofs
Appendix B Appendix - configuration and hyper-parameters
@@
For $10d$-Gaussian tasks, we use $2000$ samples per task, and a total of $T=100$ tasks.
@@
Method Discount BWT Discount BWT bound Discount forget. Discount forget. bound
@@
MLE $0.6\pm 0.0$ $2.3\pm 0.1$ $-0.3\pm 0.0$ $1.3\pm 0.1$
PAC-Bayes $0.4\pm 0.0$ $2.0\pm 0.1$ $-0.3\pm 0.1$ $1.3\pm 0.1$
Orthogonal shift
MLE $1.6\pm 0.6$ $3.2\pm 0.6$ $-0.5\pm 0.1$ $1.1\pm 0.2$
PAC-Bayes $1.3\pm 0.6$ $2.9\pm 0.5$ $-0.6\pm 0.2$ $1.0\pm 0.3$
The model consists of a shared fully connected layer of $64$ neurons and a tanh activation, as well as a linear classification head for each task. Each task was trained for $1$ epochs with batch size $64$. The learning rate was static at $1e^{-3}$. The $\lambda$ parameter was set to $1e^{-2}$. All results were run for $5$ random seeds and averages and standard error were reported.
@@
Method BWT BWT bound (3.4) Forgetting Forgetting bound (3.5) Test loss
MLE $21.1\pm 0.1$ $23.4\pm 0.4$ $-6.9\pm 0.2$ $-4.7\pm 0.4$ $12.4\pm 0.5$
PAC-Bayes $21.2\pm 0.2$ $22.8\pm 0.4$ $-6.9\pm 0.2$ $-5.2\pm 0.3$ $12.9\pm 0.3$
For split-MNIST, each task involved half of the labels (at random) chosen as positive and half as negative. This is a minor departure from the standard split-MNIST problem where $5$ different binary classification tasks are created and their loss is averaged, but the overall behavior is similar. All $60000$ training samples were used for training with a batch size of $256$. The learning rate was static at $1e^{-3}$ and the $\lambda$ parameter was set to $1e^{-4}$. The model used is identical to the CNN used for permuted-MNIST. All results were run for $5$ random seeds and averages and standard error were reported in Table 4.
@@
Method BWT BWT bound (3.4) Forgetting Forgetting bound (3.5) Test loss
MLE $19.0\pm 1.4$ $21.8\pm 1.2$ $-10.4\pm 0.4$ $-7.6\pm 0.6$ $12.8\pm 2.4$
PAC-Bayes $18.4\pm 1.1$ $21.9\pm 1.2$ $-10.2\pm 0.4$ $-6.8\pm 0.6$ $11.6\pm 1.9$
All experiments were run on local hardware with an NVIDIA GeForce 1080 GPU and a quad-core Intel i7 CPU.

### v3

@@ Appendix B Appendix - proofs
Theorem B.10.
[…]
$\begin{split}\lim_{m\rightarrow\infty}\frac{1}{T-1}\mathbb{E}_{S_{1},\ldots,S_{T}}\sum_{i=1}^{T-1}\mathcal{L}(\hat{Q}^{\sqrt{m}}_{1:T},\mathcal{D}_{i})&\leq\frac{1}{T-1}\sum_{i=1}^{T-1}\mathcal{L}(h^{*}_{i},\mathcal{D}_{i})+\lim_{m\rightarrow\infty}\frac{\mathbb{E}_{S}\sum_{i=1}^{T-1}\log\det\mathcal{L}_{i}^{\prime\prime}(h^{*}_{i})}{2\sqrt{m}(T-1)}.\end{split}$ (22)
Proof.
Starting from (19), we can use the fact that $\mathrm{cov}_{P}(i,[T])\geq 0$ to upper bound
$\begin{split}&\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})\leq\frac{\lambda_{T}K^{2}}{8m_{i}}\\ &+\frac{1}{\lambda_{T}}\mathbb{E}_{S_{T}}\log\frac{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]}{\mathbb{E}_{h\sim P}\left[e^{-\sum_{j=1,j\neq i}^{T}\lambda_{j}\hat{\mathcal{L}}(h,S_{j})}\right]\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{i})}\right]+0}\end{split}$
giving us the simplified
$\begin{split}&\mathbb{E}_{S_{i}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})\leq\frac{\lambda_{T}K^{2}}{8m_{i}}\\ &+\frac{1}{\lambda_{T}}\mathbb{E}_{S_{T}}\log\frac{1}{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\hat{\mathcal{L}}(h,S_{i})}\right]}\end{split}$
Similarly to Theorem 4.3, we can apply Jensen’s inequality and Hoeffding’s lemma, yielding
$\begin{split}&\mathbb{E}_{S}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})\leq\frac{\lambda_{T}K^{2}}{4m_{i}}\\ &+\frac{1}{\lambda_{T}}\mathbb{E}_{S_{T}}\log\frac{1}{\mathbb{E}_{h\sim P}\left[e^{-\lambda_{T}\mathcal{L}(h,\mathcal{D}_{i})}\right]}\end{split}$
Averaging out these terms, we can (under Assumption 2) apply Laplace’s method on the denominator assuming $\lambda_{T}=\sqrt{m}\rightarrow\infty$, meaning
$\begin{split}&\lim_{m\rightarrow\infty}\frac{1}{T-1}\mathbb{E}_{S}\sum_{i=1}^{T-1}\mathcal{L}(\hat{Q}^{\sqrt{m}}_{1:T},\mathcal{D}_{i})\leq\lim_{m\rightarrow\infty}\frac{1}{\sqrt{m}(T-1)}\mathbb{E}_{S}\sum_{i=1}^{T-1}\log\frac{1}{\left(\frac{2\pi}{\sqrt{m}}\right)^{d/2}\frac{P(h^{*}_{i})}{\sqrt{\det\mathcal{L}_{i}^{\prime\prime}(h^{*}_{i})}}\left[e^{-\sqrt{m}\mathcal{L}(h^{*}_{i},\mathcal{D}_{i})}\right]}\end{split}$
Considering the logarithm we can arrive at
$\begin{split}&\lim_{m\rightarrow\infty}\frac{1}{T-1}\mathbb{E}_{S}\sum_{i=1}^{T-1}\mathcal{L}(\hat{Q}^{\sqrt{m}}_{1:T},\mathcal{D}_{i})\leq\lim_{m\rightarrow\infty}\frac{1}{\sqrt{m}(T-1)}\mathbb{E}_{S}\sum_{i=1}^{T-1}\log\left(\left(\frac{2\pi}{\sqrt{m}}\right)^{-d/2}\frac{\sqrt{\det\mathcal{L}_{i}^{\prime\prime}(h^{*}_{i})}}{P(h^{*}_{i})}\left[e^{\sqrt{m}\mathcal{L}(h^{*}_{i},\mathcal{D}_{i})}\right]\right)\end{split}$
Since $m\rightarrow\infty$, we have
$\begin{split}&\lim_{m\rightarrow\infty}\frac{1}{T-1}\mathbb{E}_{S}\sum_{i=1}^{T-1}\mathcal{L}(\hat{Q}^{\sqrt{m}}_{1:T},\mathcal{D}_{i})\leq\lim_{m\rightarrow\infty}\frac{\mathbb{E}_{S}\sum_{i=1}^{T-1}\log\det\mathcal{L}_{i}^{\prime\prime}(h^{*}_{i})}{2\sqrt{m}(T-1)}+\frac{1}{(T-1)}\sum_{i=1}^{T-1}\mathcal{L}(h^{*}_{i},\mathcal{D}_{i})\end{split}$
Appendix C Appendix - configuration and hyper-parameters
@@
For $10d$-Gaussian tasks, we use $3000$ samples per task, and a total of $T=100$ tasks.
@@
Method BWT BWT bound Forgetting Forgetting bound
@@
EWC $0.4\pm 0.0$ $4\pm 0.0 0.0$$0\pm 0.0$ $0.1\pm 0.0$
VI $0.0\pm 0.0$ $1\pm 0.0 0.0$$\pm 0.0 0.0$$0\pm 0.0$
Orth. shift
EWC $1.0\pm 0.2$ $1.0\pm 0.2$ $0.1\pm 0.0$ $0.1\pm 0.0$
VI $0.3\pm 0.1$ $3\pm 0.1 0.0$$0\pm 0.0$ $0.1\pm 0.0$
The model consists of a shared fully connected layer of $64$ neurons and a tanh activation, as well as a linear classification head for each task. Each task was trained for $1$ epochs with batch size $16$. The learning rate was static at $1e^{-3}$. The $\lambda$ parameter was set to $1e^{-3}$. All results were run for $5$ random seeds and averages and standard error were reported.
For EWC, the $\sigma^{2}$ parameter used for posterior construction was set as $1e^{-2}$, and the regularization weight was set at $\lambda_{\mathrm{EWC}}=40$.
@@
Method BWT BWT bound Forgetting Forgetting bound (3.2) Test loss
EWC $1.7\pm 0.1$ $2.4\pm 0.2$ $0.6\pm 0.1$ $1.3\pm 0.2$ $1.0\pm 0.1$
VI $10.4\pm 0.3$ $12.1\pm 0.4$ $2.9\pm 0.2$ $1.2\pm 0.3$ $3.1\pm 0.1$
For split-MNIST, each task involved half of the labels (at random) chosen as positive and half as negative. This is a minor departure from the standard split-MNIST problem where $5$ different binary classification tasks are created and their loss is averaged, but the overall behavior is similar. All $60000$ training samples were used for training with a batch size of $128$. The learning rate was static at $1e^{-3}$ and the $\lambda$ parameter was set to $1e^{-3}$. The model used is identical to the CNN used for permuted-MNIST. All results were run for $5$ random seeds and averages and standard error were reported in Table 5.
@@
Method BWT BWT bound Forgetting Forgetting bound (3.2) Test loss
EWC $12.7\pm 0.7$ $3.6\pm 0.7 11$$1.9\pm 0.7$ $12.8\pm 0.7$ $0.7\pm 0.1$
VI $15.1\pm 0.3$ $17.7\pm 0.4$ $2.5\pm 0.4$ $5.1\pm 0.5$ $3.1\pm 0.3$
For split-CIFAR10, each task involved half of the labels (at random) chosen as positive and half as negative. This is a minor departure from the standard split-CIFAR10 problem where $5$ different binary classification tasks are created and their loss is averaged, but the overall behavior is similar. All $50000$ training samples were used for training with a batch size of $256$. The learning rate was static at $1e^{-3}$ and the $\lambda$ parameter was set to $1e^{-3}$. The model used is identical to the CNN used for permuted-MNIST but with $800$ neurons in the linear layer. All results were run for $5$ random seeds and averages and standard error were reported in Table 6.
Table 6: Metrics for split-CIFAR10 tasks. All test metrics are in loss percentage over the relevant models and test sets averaged over $5$ seeds. Backward transfer is the final loss percentage over all test sets. Standard error is reported after the $\pm$ sign.
Method BWT BWT bound Forgetting Forgetting bound (3.2) Test loss
EWC $36.9\pm 0.4$ $40.7\pm 0.5$ $5.7\pm 0.3$ $9.5\pm 0.6$ $30.9\pm 0.9$
VI $48.3\pm 0.0$ $50.3\pm 0.2$ $1.2\pm 0.1$ $0.8\pm 0.3$ $49.1\pm 0.4$
All experiments were run on local hardware with an NVIDIA GeForce 4090 GPU and an Intel i9 CPU.